# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v6

@@ 8.3 Scene and Demonstration Generation
Simulation scenes can be created manually. However, this approach is time-consuming and labor-intensive. As a result, automated or semi-automated scene generation methods are more preferred (Deitke et al. (2022)). Two methods can be used. Real-to-Sim method converts real scenes to simulation. Automated generation method automatically generates simulation scenes without real-world observation. Real-to-Sim method can accurately mimic the real world, but it limits the diversity of scenes. The automated generation method can create more diverse scenes and increase the variety of collected demonstrations.
@@
The Real-to-Sim method directly refers to a digital twin. The Real-to-Sim method utilizes 3D-reconstruction technology (in Sec. 6) or inverse graphics (Chen et al. (2024c)) to create the real-world scene in a virtual environment (Torne et al. (2024)). But, 3D reconstruct scene is static environment where objects lack real-world physical properties, such as material, mass and friction coefficients, and are non-interactive (Torne et al. (2024)). The inverse graphics method, such as URDFormer (Chen et al. (2024c)), directly generates interactive simulation environment and articulated objects from input RGB image. Compared to 3D-reconstruction methods, it reduces human involvement and produces interactive simulation environment. However, it lacks physical plausibility and fails to address the mismatch between the generated object’s physical properties and the real world.
@@
As for the application of foundation models in Real-to-Sim methods, GRS (Zook et al. (2024)) employs SAM2 for object segmentation from RGBD image and utilizes VLMs to describe and match objects with simulation-ready assets. This approach combines the strengths of 3D-reconstruction and inverse graphics methods. It ensures the credibility of 3D-reconstruction methods and allowing objects in the scene to interact. However, it is impossible for the assert dataset to fully cover objects in the real world. Constructing an interact assert dataset often requires manual design by the creator or human-assisted interactive object generation. ACDC (Dai et al. (2024)) defines a digital cousin concept. Unlike a digital twin, it does not directly replicate a real-world counterpart. However, it retains similar geometric and semantic features by using similar asserts when the assert dataset does not include real-world objects. As for object pose, depth cameras are commonly used, but they struggle to capture reflective surfaces accurately. This limits their use in the wild. To address this issue, ACDC uses Depth-Anything-v2 (Yang et al. (2024b)), a state-of-the-art monocular depth estimation model, to estimate the depth map.
@@
Scene diversity primarily includes the diversity of scene layouts, such as floor plans and object placements, as well as the diversity of objects. The automated generation methods are more effective for producing large-scale diverse scenes. The automated generation methods can be categorized into rule-based and learning-based approaches. For instance, ProcTHOR (Deitke et al. (2022)) introduces a procedural generation pipeline for interactive scenes using rule-based constraints and statistical priors. However, the generated scenes often rely on pre-defined priors, resulting in unrealistic outcomes that hinder agent learning (Khanna et al. (2024)). To address this, PHYSCENE (Yang et al. (2024c)) incorporates physical collision avoidance, object layouts, interactivity, and reachability metrics into a diffusion model. This approach enhances the physical plausibility and interactivity of generated scenes.
@@
Due to the prior knowledge of foundation models, there are current efforts to use foundation models for scene construction. RoboGen (Wang et al. (2023e)) utilizes LLM to generate relevant assets, asset sizes, asset configuration, scene configuration based on the task proposals and use text-to-image-to-3D generation to create the corresponding assets. These assets are imported into the simulator to generate the appropriate scene. Finally, using VLM for task-specific scene verification. GenSim (Wang et al. (2023c)) uses LLMs to generate new task and task scenario codes based on the pre-cached scene codes in a task library. However, using foundation models to automate the generation of scene’s physical plausibility still relies on VLM for judgment. At the same time, the above research also uses LLMs to generate diverse instructions to ensure task diversity. However, generating diverse task instructions with LLMs presents challenges in ensuring rationality for the current environment.
@@
The Real-to-Sim method and the Automated generation method both rely on 3D assets. The diversity of 3D assets determines the variety of scenes (Nasiriany et al. (2024)). Although there are many existing 3D object assets (Chang et al. (2015); Deitke et al. (2023); Li et al. (2023a); Geng et al. (2023a); Xiang et al. (2020); Liu et al. (2022b); Calli et al. (2017)), their quantity is far from sufficient to cover the variety of real-world objects. As a result, many studies focus on the automatic generation of assets, such as zero-1-to-3 (Wang et al. (2023e)), Luma.ai (Nasiriany et al. (2024)), LLaMA-Mesh (Wang et al. (2024e)), Trellis (Xiang et al. (2024)). However, the performance of generative models is also limited by the shortage of current 3D training data. To address this issue, data cleaning techniques or manual supervision are needed to filter and select high-quality generated object assets.
@@
The modeling of the interaction environment above primarily focuses on articulated object modeling. Articulated objects can be created manually by designers or generated using procedural (Jiang et al. (2022); Liu et al. (2023b); Zhang et al. (2023e)) or human-assisted interactive methods (Torne et al. (2024)) after 3D scanning. They can also be generated automatically through inverse graphics (Chen et al. (2024c)) or generative model (Xiang et al. (2024)). However, current automated methods for generating articulated object assets are limited to objects with few rotational joints. Real2Code (Mandi et al. (2024)) fine-tunes a CodeLlama model to process visual observation descriptions and then outputs joint predictions. This enables Real2Code to reconstruct complex articulated objects with up to 10 parts. At the same time, generative models mainly focus on rigid and articulated objects and research on deformable objects remains insufficient (Sundaresan et al. (2022)).
@@
To collect demonstrations in simulations, different approaches can be used based on task complexity. For simple tasks, like a two-finger gripper picking up a cube, a hard-coding method (Wang et al. (2022b)) can be used. However, for more complex tasks, remote teleoperation (Chen et al. (2024a)) or skill library (Ha et al. (2023)) should be employed. Building skill library can be done using reinforcement learning or gradient optimization methods. RoboGen (Wang et al. (2023e)) shows that gradient-based trajectory optimization is better for fine-grained manipulation tasks with soft bodies, like shaping dough into a specific form. On the other hand, reinforcement learning and evolutionary strategies are more effective for contact-rich tasks and continuous interactions with other components in the scene.
@@ 8.4 Sim-to-Real Gap Solutions
The sim-to-real problem is a widespread issue across machine learning, not limited to manipulation (Zhao et al. (2020)). The goal is to successfully transfer the policy from the simulation (source domain) to the real world (target domain). The gap in the manipulation tasks between the simulation and the real-world includes two main types: visual gap and dynamic gap. Visual gap refers to the difference between the vision information produced by the renderer and the vision information in the real world. In terms of rendering realism, BEHAVIOR-1K (Li et al. (2023a)) highlights that Omniverse offers the highest rendering performance. The dynamic gap consists of several factors. First, there is a difference between the physics engine used in simulations and real-world physics. Second, the properties of objects, including robots, contribute to the object dynamic gap. Lastly, there is a control gap in robots, such as variations in static errors caused by different PID parameters. Currently, there are three main approaches to address sim-to-real gap: system identification, domain randomization, and transfer learning (Zhao et al. (2020)).
@@
Most of the system identification research (Kristinsson and Dumont (1992)) aims to create an accurate mathematical model of a physical system to make the simulator more realistic. However, it is impossible to accurately build models of complex environments in simulators. The primary methods for physical parameter identification include estimation from interaction (Seker and Kroemer (2024); Bohg et al. (2017); Xu et al. (2019)), estimation from demonstrations (Torne et al. (2024)), and estimation from observations using foundation models (Gao et al. (2023)). Among these, estimation from demonstrations appears more effective. Demonstrations inherently contain interaction information and can also assist policy training. However, improving the hardware performance for collecting demonstrations remains essential.
@@
Domain randomization (Ramos et al. (2019)) involves adding random disturbances to the parameters in simulation. This can include various elements, generally divided into visual and dynamic randomization. Visual randomization covers visual parameters like lighting, object textures, and camera positions. Dynamic randomization covers dynamic parameters like object sizes, surface friction coefficients, object masses, and actuator force gains. By experiencing diverse simulated environments, the policy can adapt to a broad range of real-world conditions. For the policy, the real world is essentially just another disturbed environment. However, parameter randomization requires human expertise. Ma et al. (2024) demonstrates that LLM excels in selecting randomized parameters and determining the randomization distribution. This makes domain randomization more automated.
@@
Transfer learning (Yu and Wang (2022); Tan et al. (2018)) involves using limited real-world data to adapt a policy trained on a abundant simulation data to the real world. Treat policies in the real-world and in the simulation as different tasks. We can use task transfer methods for transfer learning. For example, Rusu et al. (2017) uses the progressive network to apply knowledge from a policy trained in simulation to a new policy trained with limited real-world data, without losing the previous knowledge. Treat the policies in the real-world and in the simulation as the same task, even though the data distributions differ. We can use domain adaptation methods to address this issue. Three common methods for domain adaptation are discrepancy-based (Lyu et al. (2024)), adversarial-based (Eysenbach et al. (2020)), and reconstruction-based methods (Bousmalis et al. (2016)). Discrepancy-based methods measure the feature distance between the source and target domains using predefined statistical metrics. This helps to align their feature spaces. Adversarial-based methods use a domain classifier to determine whether features come from the source or target domain. Once trained, the extractor can produce features that are invariant across both domains. Reconstruction-based methods also aim to find shared features between domains through setting up an auxiliary reconstruction task and using the shared features to recover the original input.
@@
The methods discussed above assume that the target domain remains unchanged. However, many physical parameters of the same robot can change significantly. Factors like temperature, humidity, positioning, and wear and tear over time can all affect these parameters. This makes it harder to bridge the sim-to-real gap. To address this issue, DORA (Zhang et al. (2024c)) uses an information bottleneck principle. It aims to maximize the mutual information between the dynamics encoding and environmental data. At the same time, it minimizes the mutual information between the dynamics encoding and the behavior policy actions. Transic (Jiang et al. (2024a)) proposes a data-driven approach that enables successful sim-to-real transfer using a human-in-the-loop framework.
@@ 8.5 Internet-Scale Video Dataset
Extensive and diverse video datasets are available from online repositories. The collection process requires querying and searching for videos with relevant content. After that, low-quality video data is removed through data cleansing. However, the raw video data cannot be directly transferred into the manipulation model due to the absence of (1) action or reward labels; (2) distribution shifts including physical embodiments, camera viewpoints, and environments. Although AVID (Smith et al. (2019)) and LbW (Xiong et al. (2021)) translate human action images from videos into robot action images, this type of translation remains limited to the pixel level; (3) essential low-level information like tactile feedback, force data, proprioceptive information, and depth perception (McCarthy et al. (2024)). However, these raw videos contain extensive visual information, such as objects, spatial information, human activities, and sequences of interactions between humans and objects (Eze and Crick (2024)). At the same time, language annotations are essential to support learning of semantic features in this visual information.
@@
Figure 8: Foundation Models for Manipulation Data Generation. Current mainstream simulators include Pybullet, MuJoCo, CoppeliaSim, NVIDIA Omniverse, and Unity. Meanwhile, learning-based generative models used as simulators have shown potential. Simulation environment generation can be classified into Real-to-Sim and Automated Generation methods. In the Real-to-Sim methods, assuming the object’s position is known, the main challenge lies in constructing the object’s 3D mesh. This can be achieved through scanning technique or by using VGM to generate the 3D mesh directly from RGB image (Chen et al. (2024c)). Additionally, GRS (Zook et al. (2024)) utilizes VLM to extract 3D object meshes corresponding to real-world object from assert database based on RGB image. In the Automated Generation methods, LLM can output scene descriptions or scene code based on task instruction. When the output is a scene description, VGM generates the objects and arranges them according to the description. Meanwhile, the generated scene need to be evaluated by VLM. (Wang et al. (2023e)). When the output is scene code, it directly generates the corresponding scene (Wang et al. (2023c)). However, this requires substantial prior knowledge of scene code within the task library. There are three methods for generating demonstrations in a scene: Hard-code, Teleoperation, and Skill Library. When building skill library, gradient optimization is effective in training skill for deformable tasks and reinforcement learning works better for contact-rich tasks (Wang et al. (2023e)). Solutions for the Sim-to-Real gap include System Identification, Domain Randomization, and Transfer Learning. For data augmentation, VFM is used to segment images first, and then VGM renders the object’s texture on the masked image.
@@
Methods to obtain language annotations are divided into manual and automated captions. Manual captions are created by humans labeling video content. Automated captions include four types: (1) Automatic Speech Recognition (ASR), which converts audio in videos to text (Xue et al. (2022)). (2) Alt-text, which collects captions from HTML alt-text attributes in web images and videos, like descriptions, tags, and titles (Bain et al. (2021)). (3) Transfer, which starts with a set of image-caption pairs. Then, captions are matched to video clips with similar frames (Nagrani et al. (2022)). (4) Foundation Models, which use pre-trained models to get captions. For example, VLMs provide single-frame image captions, while LLMs filter out inconsistent captions across frames (Blattmann et al. (2023)). Owing to recent advancements in language annotation techniques, most widely used internet video datasets incorporate language annotations, such as InternVid (Wang et al. (2023d)), HD-VILA-100M (Xue et al. (2022)), YT-Temporal-180M (Zellers et al. (2021)), WTS-70M (Shvetsova et al. (2025)), HowTo100M (Miech et al. (2019)), WebVid-10M (Nan et al. (2024)), and VideoCC3M (Yan et al. (2022)). At the same time, various off-the-shelf models can be used to annotate the current video with additional labels, such as pose (Shaw et al. (2023)), affordance (Mendonca et al. (2023)), key points trajectory (Wen et al. (2023b)), latent action (Ye et al. (2024)), mask and bounding boxes (Shan et al. (2020)).
@@
The task information contained in internet video data may not be highly relevant to the specific tasks performed by robots. Additionally, internet video data often suffers from issues such as missing action labels, low-level information, and distribution shifts. Therefore, manually recording custom videos can be an effective approach to collecting videos that are directly relevant to specific robot tasks or embodiments. This method can also help avoid the issue of re-annotating. By incorporating sensors such as IMUs, tactile sensors, and depth sensors during the recording process, manually recorded custom videos can exhibit lower noise compared to internet video data. However, the scale and diversity of manually recorded videos still cannot match the internet video data (McCarthy et al. (2024)). Currently, there are several commonly used manually recorded video datasets, such as Ego-4D (Grauman et al. (2022)), Ego-Exo-4D (Grauman et al. (2024)), RoboVQA (Sermanet et al. (2024)), Epic-Kitchens-100 (Damen et al. (2022)) and ActionSense (DelPreto et al. (2022)).

### v7

@@ 8.3 Scene and Demonstration Generation
Simulation scenes can be created manually. However, this approach is time-consuming and labor-intensive. As a result, automated or semi-automated scene generation methods are more preferred (deitke202). Two methods can be used. Real-to-Sim method converts real scenes to simulation. Automated generation method automatically generates simulation scenes without real-world observation. Real-to-Sim method can accurately mimic the real world, but it limits the diversity of scenes. The automated generation method can create more diverse scenes and increase the variety of collected demonstrations.
@@
The Real-to-Sim method directly refers to a digital twin. The Real-to-Sim method utilizes 3D-reconstruction technology (in Sec. 6) or inverse graphics (chen2024urdformer) to create the real-world scene in a virtual environment (torne2024reconciling). But, 3D reconstruct scene is static environment where objects lack real-world physical properties, such as material, mass and friction coefficients, and are non-interactive (torne2024reconciling). The inverse graphics method, such as URDFormer (chen2024urdformer), directly generates interactive simulation environment and articulated objects from input RGB image. Compared to 3D-reconstruction methods, it reduces human involvement and produces interactive simulation environment. However, it lacks physical plausibility and fails to address the mismatch between the generated object’s physical properties and the real world.
@@
As for the application of foundation models in Real-to-Sim methods, GRS (zook2024grs) employs SAM2 for object segmentation from RGBD image and utilizes VLMs to describe and match objects with simulation-ready assets. This approach combines the strengths of 3D-reconstruction and inverse graphics methods. It ensures the credibility of 3D-reconstruction methods and allowing objects in the scene to interact. However, it is impossible for the assert dataset to fully cover objects in the real world. Constructing an interact assert dataset often requires manual design by the creator or human-assisted interactive object generation. ACDC (dai2024automated) defines a digital cousin concept. Unlike a digital twin, it does not directly replicate a real-world counterpart. However, it retains similar geometric and semantic features by using similar asserts when the assert dataset does not include real-world objects. As for object pose, depth cameras are commonly used, but they struggle to capture reflective surfaces accurately. This limits their use in the wild. To address this issue, ACDC uses Depth-Anything-v2 (yang2024depth), a state-of-the-art monocular depth estimation model, to estimate the depth map.
@@
Scene diversity primarily includes the diversity of scene layouts, such as floor plans and object placements, as well as the diversity of objects. The automated generation methods are more effective for producing large-scale diverse scenes. The automated generation methods can be categorized into rule-based and learning-based approaches. For instance, ProcTHOR (deitke202) introduces a procedural generation pipeline for interactive scenes using rule-based constraints and statistical priors. However, the generated scenes often rely on pre-defined priors, resulting in unrealistic outcomes that hinder agent learning (khanna2024habitat). To address this, PHYSCENE (yang2024physcene) incorporates physical collision avoidance, object layouts, interactivity, and reachability metrics into a diffusion model. This approach enhances the physical plausibility and interactivity of generated scenes.
@@
Due to the prior knowledge of foundation models, there are current efforts to use foundation models for scene construction. RoboGen (wang2023robogen) utilizes LLM to generate relevant assets, asset sizes, asset configuration, scene configuration based on the task proposals and use text-to-image-to-3D generation to create the corresponding assets. These assets are imported into the simulator to generate the appropriate scene. Finally, using VLM for task-specific scene verification. GenSim (wang2023gensim) uses LLMs to generate new task and task scenario codes based on the pre-cached scene codes in a task library. However, using foundation models to automate the generation of scene’s physical plausibility still relies on VLM for judgment. At the same time, the above research also uses LLMs to generate diverse instructions to ensure task diversity. However, generating diverse task instructions with LLMs presents challenges in ensuring rationality for the current environment.
@@
The Real-to-Sim method and the Automated generation method both rely on 3D assets. The diversity of 3D assets determines the variety of scenes (nasiriany2024robocasa). Although there are many existing 3D object assets (chang2015shapenet; deitke2023objaverse; li2023behavior; geng2023gapartnet; xiang2020sapien; liu2022akb; calli2017yale), their quantity is far from sufficient to cover the variety of real-world objects. As a result, many studies focus on the automatic generation of assets, such as zero-1-to-3 (wang2023robogen), Luma.ai (nasiriany2024robocasa), LLaMA-Mesh (wang2024llama), Trellis (xiang2024structured). However, the performance of generative models is also limited by the shortage of current 3D training data. To address this issue, data cleaning techniques or manual supervision are needed to filter and select high-quality generated object assets.
@@
The modeling of the interaction environment above primarily focuses on articulated object modeling. Articulated objects can be created manually by designers or generated using procedural (jiang2022ditto; liu2023paris; zhang2023part) or human-assisted interactive methods (torne2024reconciling) after 3D scanning. They can also be generated automatically through inverse graphics (chen2024urdformer) or generative model (xiang2024structured). However, current automated methods for generating articulated object assets are limited to objects with few rotational joints. Real2Code (mandi2024real2code) fine-tunes a CodeLlama model to process visual observation descriptions and then outputs joint predictions. This enables Real2Code to reconstruct complex articulated objects with up to 10 parts. At the same time, generative models mainly focus on rigid and articulated objects and research on deformable objects remains insufficient (sundaresan2022diffcloud).
@@
To collect demonstrations in simulations, different approaches can be used based on task complexity. For simple tasks, like a two-finger gripper picking up a cube, a hard-coding method (wang2022bulletarm) can be used. However, for more complex tasks, remote teleoperation (chen2024teleoperation) or skill library (ha2023scaling) should be employed. Building skill library can be done using reinforcement learning or gradient optimization methods. RoboGen (wang2023robogen) shows that gradient-based trajectory optimization is better for fine-grained manipulation tasks with soft bodies, like shaping dough into a specific form. On the other hand, reinforcement learning and evolutionary strategies are more effective for contact-rich tasks and continuous interactions with other components in the scene.
@@ 8.4 Sim-to-Real Gap Solutions
The sim-to-real problem is a widespread issue across machine learning, not limited to manipulation (zhao2020sim). The goal is to successfully transfer the policy from the simulation (source domain) to the real world (target domain). The gap in the manipulation tasks between the simulation and the real-world includes two main types: visual gap and dynamic gap. Visual gap refers to the difference between the vision information produced by the renderer and the vision information in the real world. In terms of rendering realism, BEHAVIOR-1K (li2023behavior) highlights that Omniverse offers the highest rendering performance. The dynamic gap consists of several factors. First, there is a difference between the physics engine used in simulations and real-world physics. Second, the properties of objects, including robots, contribute to the object dynamic gap. Lastly, there is a control gap in robots, such as variations in static errors caused by different PID parameters. Currently, there are three main approaches to address sim-to-real gap: system identification, domain randomization, and transfer learning (zhao2020sim).
@@
Most of the system identification research (kristinsson1992system) aims to create an accurate mathematical model of a physical system to make the simulator more realistic. However, it is impossible to accurately build models of complex environments in simulators. The primary methods for physical parameter identification include estimation from interaction (seker2024estimating; bohg2017interactive; xu2019densephysnet), estimation from demonstrations (torne2024reconciling), and estimation from observations using foundation models (gao2023physically). Among these, estimation from demonstrations appears more effective. Demonstrations inherently contain interaction information and can also assist policy training. However, improving the hardware performance for collecting demonstrations remains essential.
@@
Domain randomization (ramos2019bayessim) involves adding random disturbances to the parameters in simulation. This can include various elements, generally divided into visual and dynamic randomization. Visual randomization covers visual parameters like lighting, object textures, and camera positions. Dynamic randomization covers dynamic parameters like object sizes, surface friction coefficients, object masses, and actuator force gains. By experiencing diverse simulated environments, the policy can adapt to a broad range of real-world conditions. For the policy, the real world is essentially just another disturbed environment. However, parameter randomization requires human expertise. ma2024dreureka demonstrates that LLM excels in selecting randomized parameters and determining the randomization distribution. This makes domain randomization more automated.
@@
Transfer learning (yu2022dexterous; tan2018survey) involves using limited real-world data to adapt a policy trained on a abundant simulation data to the real world. Treat policies in the real-world and in the simulation as different tasks. We can use task transfer methods for transfer learning. For example, rusu2017sim uses the progressive network to apply knowledge from a policy trained in simulation to a new policy trained with limited real-world data, without losing the previous knowledge. Treat the policies in the real-world and in the simulation as the same task, even though the data distributions differ. We can use domain adaptation methods to address this issue. Three common methods for domain adaptation are discrepancy-based (lyu2024cross), adversarial-based (eysenbach2020off), and reconstruction-based methods (bousmalis2016domain). Discrepancy-based methods measure the feature distance between the source and target domains using predefined statistical metrics. This helps to align their feature spaces. Adversarial-based methods use a domain classifier to determine whether features come from the source or target domain. Once trained, the extractor can produce features that are invariant across both domains. Reconstruction-based methods also aim to find shared features between domains through setting up an auxiliary reconstruction task and using the shared features to recover the original input.
@@
The methods discussed above assume that the target domain remains unchanged. However, many physical parameters of the same robot can change significantly. Factors like temperature, humidity, positioning, and wear and tear over time can all affect these parameters. This makes it harder to bridge the sim-to-real gap. To address this issue, DORA (zhang2024debiased) uses an information bottleneck principle. It aims to maximize the mutual information between the dynamics encoding and environmental data. At the same time, it minimizes the mutual information between the dynamics encoding and the behavior policy actions. Transic (jiang2024transic) proposes a data-driven approach that enables successful sim-to-real transfer using a human-in-the-loop framework.
@@ 8.5 Internet-Scale Video Dataset
Extensive and diverse video datasets are available from online repositories. The collection process requires querying and searching for videos with relevant content. After that, low-quality video data is removed through data cleansing. However, the raw video data cannot be directly transferred into the manipulation model due to the absence of (1) action or reward labels; (2) distribution shifts including physical embodiments, camera viewpoints, and environments. Although AVID (smith2019avid) and LbW (xiong2021learning) translate human action images from videos into robot action images, this type of translation remains limited to the pixel level; (3) essential low-level information like tactile feedback, force data, proprioceptive information, and depth perception (mccarthy2024towards). However, these raw videos contain extensive visual information, such as objects, spatial information, human activities, and sequences of interactions between humans and objects (eze2024learning). At the same time, language annotations are essential to support learning of semantic features in this visual information.
@@
Figure 8: Foundation Models for Manipulation Data Generation. Current mainstream simulators include Pybullet, MuJoCo, CoppeliaSim, NVIDIA Omniverse, and Unity. Meanwhile, learning-based generative models used as simulators have shown potential. Simulation environment generation can be classified into Real-to-Sim and Automated Generation methods. In the Real-to-Sim methods, assuming the object’s position is known, the main challenge lies in constructing the object’s 3D mesh. This can be achieved through scanning technique or by using VGM to generate the 3D mesh directly from RGB image (chen2024urdformer). Additionally, GRS (zook2024grs) utilizes VLM to extract 3D object meshes corresponding to real-world object from assert database based on RGB image. In the Automated Generation methods, LLM can output scene descriptions or scene code based on task instruction. When the output is a scene description, VGM generates the objects and arranges them according to the description. Meanwhile, the generated scene need to be evaluated by VLM. (wang2023robogen). When the output is scene code, it directly generates the corresponding scene (wang2023gensim). However, this requires substantial prior knowledge of scene code within the task library. There are three methods for generating demonstrations in a scene: Hard-code, Teleoperation, and Skill Library. When building skill library, gradient optimization is effective in training skill for deformable tasks and reinforcement learning works better for contact-rich tasks (wang2023robogen). Solutions for the Sim-to-Real gap include System Identification, Domain Randomization, and Transfer Learning. For data augmentation, VFM is used to segment images first, and then VGM renders the object’s texture on the masked image.
@@
Methods to obtain language annotations are divided into manual and automated captions. Manual captions are created by humans labeling video content. Automated captions include four types: (1) Automatic Speech Recognition (ASR), which converts audio in videos to text (xue2022advancing). (2) Alt-text, which collects captions from HTML alt-text attributes in web images and videos, like descriptions, tags, and titles (bain2021frozen). (3) Transfer, which starts with a set of image-caption pairs. Then, captions are matched to video clips with similar frames (nagrani2022learning). (4) Foundation Models, which use pre-trained models to get captions. For example, VLMs provide single-frame image captions, while LLMs filter out inconsistent captions across frames (blattmann2023stable). Owing to recent advancements in language annotation techniques, most widely used internet video datasets incorporate language annotations, such as InternVid (wang2023internvid), HD-VILA-100M (xue2022advancing), YT-Temporal-180M (zellers2021merlot), WTS-70M (shvetsova2025howtocaption), HowTo100M (miech2019howto100m), WebVid-10M (nan2024openvid), and VideoCC3M (yan2022videococa). At the same time, various off-the-shelf models can be used to annotate the current video with additional labels, such as pose (shaw2023videodex), affordance (mendonca2023structured), key points trajectory (wen2023any), latent action (ye2024latent), mask and bounding boxes (shan2020understanding).
@@
The task information contained in internet video data may not be highly relevant to the specific tasks performed by robots. Additionally, internet video data often suffers from issues such as missing action labels, low-level information, and distribution shifts. Therefore, manually recording custom videos can be an effective approach to collecting videos that are directly relevant to specific robot tasks or embodiments. This method can also help avoid the issue of re-annotating. By incorporating sensors such as IMUs, tactile sensors, and depth sensors during the recording process, manually recorded custom videos can exhibit lower noise compared to internet video data. However, the scale and diversity of manually recorded videos still cannot match the internet video data (mccarthy2024towards). Currently, there are several commonly used manually recorded video datasets, such as Ego-4D (grauman2022ego4d), Ego-Exo-4D (grauman2024ego), RoboVQA (sermanet2024robovqa), Epic-Kitchens-100 (damen2022rescaling) and ActionSense (delpreto2022actionsense).